Shift report 20010910 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Sep 10 2001 - 16:07:49 EDT

  • Next message: Flemming Videbaek: "Re: Optical Algorithm for RHIC PID"

    ------------------------------------------------------------
    Date: Mon Sep 10 16:07:49 2001, Shift: 08:00-16:00
    Supervisor: Eun-Joo, with: FV, Dana
    Log book#8 pages 105 - 106
    
    Shift Summary:
    ------------------------------------------------------------
    1. No useful beam. (beam study with 6-bunch beam)
    
    2. Run 4665 : Camac Pedestal Run
       Run 4688 : TPC Pedestal Run
       Run 4689 : TDC Calibration Run for Dana
    
    3. No detector Problems.
    
       Dana fixed H1 TDC Bot33, Top15.
       DAQ problem when I tried to TPC Pedestal Run
         -> Problems : FB3-Comm error
                       Trigger 1 increased instead of Trigger 7
         -> I followed Hiro's shift report in Sep-01.
         -> I stop/start IOServer, RunServer in DAQ tasks 
            many times (more than 10).
         ->  Suddenly fixed, and I don't know why/how.     
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Sep 10 2001 - 16:08:42 EDT